# Random Long-Context Access for Mamba via Hardware-aligned Hierarchical Sparse Attention

### v1

@@ Abstract
A key advantage of Recurrent Neural Networks (RNNs) over Transformers is their linear computational and space complexity enables faster training and inference for long sequences. However, RNNs are fundamentally unable to randomly access historical context, and simply integrating attention mechanisms may undermine their efficiency advantages. To overcome this limitation, we propose Hierarchical Sparse Attention (HSA), a novel attention mechanism that enhances RNNs with long-range random access flexibility while preserving their merits in efficiency and length generalization. HSA divides inputs into chunks, selecting the top- chunks and hierarchically aggregates information. The core innovation lies in learning token-to-chunk relevance based on fine-grained token-level information inside each chunk. This approach enhances the precision of chunk selection across both in-domain and out-of-domain context lengths. To make HSA efficient, we further introduce a hardware-aligned kernel design. By combining HSA with Mamba, we introduce RAMba, which achieves perfect accuracy in passkey retrieval across 64 million contexts despite pre-training on only 4K-length contexts, and significant improvements on various downstream tasks, with nearly constant memory footprint. These results show RAMba’s huge potential in long-context modeling. 00footnotetext: ‡ Corresponding authors
@@ 1 Introduction
However, RNN-based models suffer from a critical limitation: an information bottleneck [55] caused by compressing variable-length contexts into fixed-dimensional representations. Unlike attention mechanisms, they lack random access to contextual information, which becomes especially problematic in tasks like passkey retrieval, where performance degrades as sequence length increases [59]. While augmenting RNNs with attention mechanisms can help mitigate this limitation, it introduces new challenges—such as poor length extrapolation, quadratic computational complexity, and substantial memory overhead during inference—ultimately undermining the original efficiency advantages of RNNs. Consequently, there remains no satisfactory RNN-based solution that can simultaneously achieve length generalization, random-access flexibility, and efficiency.
@@
To address the trilemma, we propose a novel Hierarchical Sparse Attention (HSA) mechanism. Existing sparse attention methods [66, 31] typically divide sequences into chunks, allowing each token to attend to a concatenation of selected chunks, achieving efficiency while retraining random access capability. While this design offers a promising step toward resolving the trilemma, a closer examination reveals a critical weakness: these methods often suffer from inaccurate chunk selection, both within the training distribution and when generalizing to longer, out-of-domain contexts. As illustrated in Figure 1(A), the issue may stem from estimating token-to-chunk relevance via attention scores, which are learned from token-to-token gradients instead of feedback from entire chunks.
@@
Based on this insight, HSA introduces a two-stage hierarchical mechanism to enable end-to-end learning of token-to-chunk relevance. As shown in Figure 1(B), in the first stage, it separately weights over tokens within each chunk to capture chunk-level information. In the second stage, it fuses the chunk information with weights derived from token-to-chunk scores. During the backpropagation pass, the chunks that are more helpful for next-token prediction are assigned higher weights. Experimental results show that this chunk selection mechanism remains accurate even when context lengths exceed pre-training lengths by 1000 times under a certain condition. Since each token corresponds to different chunks, a naive implementation would require substantial memory. Thus, we further propose a hardware-aligned algorithm to achieve efficient parallel computation.
@@
Built upon HSA, we propose Random Access Mamba (RAMba), an extension of Mamba-2 [16] that incorporates HSA into specific layers. To maintain a constant memory footprint during inference, RAMba offloads the token-level key-value (KV) cache to CPU memory while retaining compact chunk-level representations on the GPU for efficient chunk selection. At each time step, only the KV cache of the selected chunks is loaded onto the GPU, ensuring efficient memory usage. To further minimize GPU-CPU memory swaps, RAMba leverages the hidden states derived from an intermediate layer as shared KV cache for all subsequent HSA layers, requiring just one chunk selection and swap per step. This architecture can simultaneously balance training and inference efficiency, length generalization ability, and long-range random access flexibility.
@@
In our experiments, we compare RAMba against baselines like Transformers, Mamba-2, their variants with sliding window attention [14] and NSA [66] across long-range language modeling, downstream tasks, and efficiency. RAMba consistently outperforms the baselines in long-context modeling and downstream tasks while exhibiting exceptional length generalization. Notably, it is the first Mamba-based model to achieve perfect accuracy on a 64M context in the passkey retrieval task. In terms of efficiency, HSA is three times faster than NSA and 5–25 times faster than flash attention with 16K context length or longer in forward pass. Additionally, With memory offloading enabled, RAMba maintains nearly constant memory usage. These results demonstrate RAMba’s superior capability in long-text modeling. In summary, our contributions are three fold:
@@
We propose HSA, a novel hierarchical attention mechanism with a hardware-efficient algorithm, offering efficiency, length generalization, and long-range random access flexibility.
@@ 2 Related Works
Sparse attention aims to reduce computational complexity by focusing on a limited number of tokens. For example, sliding window attention and its variants [40, 14, 67, 9] restrict computations to a fixed-size local window for each token. Such methods often sacrifice the ability to capture long-range information. Clustering based approaches [29, 58, 61, 51] employs locality-sensitive hashing or K-Means for token clustering and perform attention with clusters. These methods often trade off efficiency for quality due to the limited accuracy of clustering. Cache eviction approaches [70, 21, 30] maintain constant memory costs by retaining only the most important tokens in the KV cache, but this limits the model’s ability to access arbitrary contexts randomly. Combiner [49] utilizes a hierarchical attention mechanism by compressing tokens within chunks into single key-value pairs via max-pooling. uowever, this approach sacrifice the ability to capture token-level information inside the chunk, resulting in reduced accuracy. Recently, NSA [66] and MoBA [31] achieve sparsity by dividing a sequence into chunks and dynamically selecting relevant chunks for each token based on summed token-level attention scores. However, these approaches struggle with accurately identifying important chunks. Our work introduces a two-stage hierarchical attention mechanism with sparse chunk selection, achieving end-to-end learning of token-to-chunk relevance, while maintaining the ability to capture token-level information, resulting in attention flexibility, accuracy, and efficiency.
@@
RNN-based language models have recently gained increased attention because their per-token inference cost remains constant as the sequence length grows. Linear Attention [27] shows that replacing softmax attention with kernel-based approximations allows Transformers to be reformulated as RNNs, achieving similar performance while benefiting from recurrent properties. Recent advancements in RNNs, such as RWKV series [41, 42], Mamba series [23, 17], GLA [63], HGRN series [45, 46], xLSTM [8] and others [69, 64], continue to push the boundaries of this approach. Our work builds on these RNN models and is in principle applicable across different RNN architectures. Meanwhile, other studies [18, 36, 50] explore hybrid architectures of RNNs and attention by leveraging for long-range memory and attention for short-range patterns. In contrast, our approach leverages RNNs to adaptively retain essential short-range information for next-token prediction, while sparse attention selectively accesses long-range context at arbitrary positions.
@@ 3.1 Model Architecture
As shown in Figure 2(a), RAMba contains Mamba layers, equally divided into upper and lower decoders. Given the input sequence , their corresponding embeddings are fed into the lower decoder and the outputs are divided into chunks according to chunk size . These chunked representations are then passed through a Transformer-based bi-directional encoder independently to form chunked memories, used for chunk selection and sparse attention in subsequent HSA layers. The upper decoder takes the outputs of the lower decoder as inputs, in which HSA layers alternate with Mamba layers, repeating times. HSA layers capture long-range dependencies by attending to the selected chunks. Between the lower and upper decoders, a chunk selection layer selects the most relevant chunks based on the dot-product similarity between token and chunk representations.
@@
We focus on architectures like GQA [2], where every query heads shares a common KV head. The hidden size is divided into groups, satisfying , with . Here, is the dimension of each group, and is the dimension per head. Since all groups perform chunk selection and HSA identically, subsequent notation refers to an arbitrary group representative of all. Specifically, denotes one group of the output of the -th layer at step , and denotes one group of the encoded token representations of the -th chunk.
@@
At each time step , each token uses the representation derived from the lower decoder to select top- past chunks. Each query group independently selects these chunks, which are shared across the heads within the group. Formally, for each group, we have:
where is RMSNorm [68], denotes the ranking position in descending order, is the mean-pooled representations within the -th chunk, and represents token and chunk representations for chunk selections, is the relevance score of to the -th chunk, and is the selected chunk indices for .
@@ 3.1.1 Hierarchical Sparse Attention
To avoid the impact of position encoding on length generalization [28], we discard it and instead model the ordinal relationships of distances akin to the stick-breaking attention [53]. In the stick-breaking process, participants sequentially take a portion from a stick, with later participants dividing the remaining portion left by earlier ones, thus introducing a “most recent” bias. Analogously, let the total weights serve as the stick and denote the weight assigned to the -th chunk selected by the -th token, we have:
@@
where are the key and value corresponding to the -th chunk, which are shared across all HSA layers. is the query representation of at the -th layer and is obtained by applying RMSNorm [68] followed by a linear transformation.
@@ 3.1.2 Kernel Design
In HSA, each token corresponds to a distinct group of chunks, which can lead to a substantial memory footprint in a naive implementation. Inspired by NSA, we address this issue by implementing hardware-aligned HSA kernels based on Triton [54]. As illustrated in Figure 2(b), each GPU thread loads the query representations for a single step along with the selected chunks’ KV pairs for attention computation. The Algorithm 1 demonstrates a parallel forward pass for one group of query heads, with multiple groups processed in parallel following the same approach. Particularly, we use softmax-off-by-one [33] to allow tokens in the current chunk to ignore any retrieved tokens. During the back-propagation process, we adopt a two-phased backward pass inspired by Dao [15]. The first stage accumulates gradients for and , followed by and in the second stage. The pseudo-code for this process is shown in Algorithm 2 and 3, in which denotes whether the -th token selectes the -th chunk, and represents the index of the -th chunk in
@@ 3.2 Training & Inference
Many efforts have been made to address Mamba’s length generalization issues. A simple yet effective solution is truncated backpropagation through time (BPTT) [13, 63], in which the first state of a sequence is initialized as the final state of its preceding sequence. We follow this approach in training RAMba and other baselines. However, even though RAMba demonstrates certain extrapolation capabilities, we still observe an increase in perplexity on longer contexts. We hypothesize that RNNs’ memory state provides certain shortcuts [22] for long-range attention, degrading performance on contexts largely exceeding the pre-trained length. Thus we attempt to introduce appropriate forgetting mechanism into the memory state to disrupt the shortcuts. A straightforward way is memory reset, where sequences are divided into equal segments, and the initial state of each segment is reset to zero. To align with BPTT, we set the initial state as the last hidden state of a random segment in the previous step. Our experiments show that this straightforward technique effectively improves length generalization for RAMba.
@@
A key challenge in introducing attention mechanisms to RNNs is managing the memory footprint, as the KV cache scales linearly with sequence length. Prior works [34, 26] have demonstrated the feasibility of offloading the KV cache to CPU memory and selectively loading chunks during inference. For a prompt of length , all chunk representations are offloaded to the CPU after prefilling, while only is kept in GPU memory for chunk selection. During decoding, RAMba retrieves and loads chunks at each step. Since the KV cache is shared across all HSA layers, the number of parameters exchanged between the CPU and GPU totals , with per group head number usually set to 16. Our efficiency analysis in § 4.4 demonstrates the overhead for memory exchange is fully acceptable. Since the memory footprint of still increase with the sequence length, to theoretically achieving constant memory, a straight forward way is to offloading it to a FAISS [19] database. However, in practice, the memory footprint of is very limited, making such offloading unnecessary. Detailed analysis is elaborated in the experiments section.
@@ 4 Experiments
To ensure a fair comparison, we pre-train all models from scratch with 4K context length to observe their performance and extrapolation capabilities across various tasks.
@@ 4.1 Setups
We adopt the Mamba-2 architecture as the backbone of the RNN model and YaRN [43] as the Transformer baseline. The parameter size of all models trained from scratch is 370M, with detailed parameters provided in Appendix A. We study Mamba variants with different attention mechanisms, including sliding window attention, native sparse attention (NSA), and HSA. For sliding window attention, the window size is set to 512, incorporating two position encoding schemes: ALibi [44] and RoPE, the latter following the settings in Samba [50]. We set the chunk size of HSA to 64 following NSA. To ensure that the field of view for sparse attention matches the sliding window size (64 * 8 = 512), we set the number of selected chunks to 8. Regarding the NSA, we use its efficient open-source implementation 111https://github.com/fla-org/native-sparse-attention. To purely compare the capabilities of the sparse attention components, we disable NSA’s sliding window attention. The HSA incorporates a single-layered Transformer-based bi-directional encoder for chunk memory encoding, accounting for 5.4% of the total parameters. As it does not participate in decoding, its impact on fairness is minimal. HSA layers are inserted into the upper decoder every Mamba layers, while the placement of other attention layers remains consistent with this pattern. These settings remain consistent across all subsequent 370M models. Since the compressed attention in NSA functions similarly to Combiner [49], we don’t conduct a separate comparison against it. Some other related works [35] are not included in the experiments due to the lack of a well-implemented open-source reproduction.
Memory Reset.
We use w/ m.r. and w/o m.r. to denote whether the memory reset is enabled during training. Specifically, for an 8K token context, the memory is reset every 4K tokens, which aligns with the context length of other baselines. However, the chunk selection scope for sparse attention spans 8K tokens, which might be unfair to other baselines. To ensure fair comparisons, we apply the same settings to both HSA and NSA.

### v2

# Hardware-aligned Hierarchical Sparse Attention for Efficient Long-term Memory Access

@@ Abstract
A key advantage of Recurrent Neural Networks (RNNs) over Transformers is their linear computational and space complexity enables faster training and inference for long sequences. However, RNNs are fundamentally unable to randomly access historical context, and simply integrating attention mechanisms may undermine their efficiency advantages. To overcome this limitation, we propose Hierarchical Sparse Attention (HSA), a novel attention mechanism that enhances RNNs with long-range random access flexibility while preserving their merits in efficiency and length generalization. HSA divides inputs into chunks, selects the top- chunks and hierarchically aggregates information. The core innovation lies in learning token-to-chunk relevance based on fine-grained token-level information inside each chunk. This approach enhances the precision of chunk selection across both in-domain and out-of-domain context lengths. To make HSA efficient, we further introduce a hardware-aligned kernel design. By combining HSA with Mamba, we introduce RAMba, which achieves perfect accuracy in passkey retrieval across 64 million contexts despite pre-training on only 4K-length contexts, and significant improvements on various downstream tasks, with nearly constant memory footprint. These results show RAMba’s huge potential in long-context modeling.
@@ 1 Introduction
However, RNN-based models suffer from a critical limitation: the information bottleneck [58] caused by compressing variable-length contexts into fixed-dimensional representations. Unlike attention mechanisms, they lack random access to contextual information, which becomes especially problematic in tasks like passkey retrieval, where performance degrades as sequence length increases [62]. While augmenting RNNs with attention mechanisms can help mitigate this limitation, it introduces drawbacks such as poor length extrapolation, quadratic computational complexity, and substantial memory overhead during inference, ultimately undermining the original efficiency advantages of RNNs. Consequently, there remains no satisfactory RNN-based solution that can simultaneously achieve length generalization, random-access flexibility, and efficiency.
@@
To address the trilemma, we propose a novel Hierarchical Sparse Attention (HSA) mechanism. Existing sparse attentions like NSA and MoBA [70, 32] typically divide sequences into chunks, allowing each token to attend to a concatenation of selected chunks, thus can potentially achieve both efficiency and random access. While this design offers a promising step toward resolving the trilemma, a closer examination reveals a critical weakness: these methods often suffer from inaccurate chunk selection, both within the training distribution and when generalizing to longer, out-of-domain contexts. As illustrated in Figure 1(A), the issue may stem from learning token-to-chunk relevance in a chunk-unaware way, relying on token-to-token gradients instead of chunk-level feedback.
@@
Based on this insight, HSA introduces a two-stage hierarchical mechanism for the selected chunks to enable end-to-end learning of token-to-chunk relevance. As shown in Figure 1(B), in the first stage, it applies attention separately to the tokens within each chunk to capture chunk-level information. In the second stage, it fuses the chunk-level information by applying weighted summation using token-to-chunk weights. During the backpropagation pass, token-to-chunk weights are adjusted based on the contribution of the entire chunk to the next token prediction, achieving chunk-aware learning. Experimental results show that this chunk selection mechanism remains accurate even when context lengths exceed pre-training lengths by over 10,000 times in passkey retrieval. Since each token corresponds to different chunks, a naive implementation would require substantial memory. Thus, we further propose a hardware-aligned algorithm to achieve efficient parallel computation.
@@
With HSA, we propose Random Access Mamba (RAMba), an extension of Mamba-2 [16] that incorporates HSA into specific layers. To maintain a constant memory footprint during inference, RAMba offloads the token-level key-value (KV) cache to CPU memory while retaining compact chunk-level representations on the GPU for efficient chunk selection. At each time step, only the KV cache of the selected chunks is loaded onto the GPU, ensuring efficient memory usage. To further minimize GPU-CPU memory swaps, RAMba leverages the hidden states derived from an intermediate layer as shared KV cache for all subsequent HSA layers, requiring just one chunk selection and swap per step. This architecture can simultaneously balance training and inference efficiency, length generalization, and long-range random access flexibility.
@@
In our experiments, we compare RAMba with baselines such as Transformers, Mamba-2, and their variants with sliding window attention [14] and NSA [70], evaluating performance across long-range language modeling, downstream tasks, and efficiency. RAMba consistently outperforms the baselines in long-context modeling and downstream tasks while exhibiting exceptional length generalization. Notably, it is the first Mamba-based model to achieve perfect accuracy on a 64M context in the passkey retrieval task. In terms of efficiency, HSA is 3 faster than NSA and 5–25 faster than full attention for contexts of 16K tokens or more during the forward pass. Additionally, when memory offloading is enabled, RAMba maintains nearly constant memory usage. These results demonstrate RAMba’s superior capability in long-text modeling. In summary, our contributions are threefold:
@@
We propose HSA, a novel hierarchical attention mechanism paired with a hardware-efficient algorithm that simultaneously enables efficiency, length generalization, and flexible long-range random access.
@@ 2 Related Works
Sparse attention aims to reduce computational complexity by focusing on a limited number of tokens. For example, sliding window attention and its variants [41, 14, 71, 9] restrict computations to a fixed-size local window for each token. Such methods often sacrifice the ability to capture long-range information. Clustering based approaches [30, 61, 64, 52] employ locality-sensitive hashing or K-Means for token clustering and perform attention with clusters. These methods often trade off efficiency for quality due to the limited accuracy of clustering. Cache eviction approaches [74, 21, 31] maintain constant memory costs by retaining only the most important tokens in the KV cache, but this limits the model’s ability to access arbitrary contexts randomly. Combiner [50] utilizes a hierarchical attention mechanism by compressing tokens within chunks into single key-value pairs via max-pooling. However, this approach sacrifices the ability to capture token-level information inside the chunk, resulting in reduced accuracy. Recently, NSA [70] and MoBA [32] achieve sparsity by dividing a sequence into chunks and dynamically selecting relevant chunks for each token based on summed token-level attention scores. However, these approaches struggle with accurately identifying important chunks. Our work introduces a two-stage hierarchical attention mechanism with sparse chunk selection, achieving end-to-end learning of token-to-chunk relevance, while maintaining the ability to capture token-level information, resulting in attention flexibility, accuracy, and efficiency.
@@
RNN-based language models have recently gained increased attention because their per-token inference cost remains constant as the sequence length grows. Linear Attention [28] shows that replacing softmax attention with kernel-based approximations allows Transformers to be reformulated as RNNs, achieving similar performance while benefiting from recurrent properties. Recent advancements in RNNs, such as RWKV series [42, 43], Mamba series [23, 17], GLA [66], HGRN series [46, 47], xLSTM [8] and others [73, 67], continue to push the boundaries of this approach. Our work builds on these RNN models and is in principle applicable across different RNN architectures. Meanwhile, other studies [18, 37, 51] explore hybrid architectures using RNNs for long-range memory and attention for short-range patterns. In contrast, our approach leverages RNNs to adaptively retain essential short-range information for next-token prediction, while sparse attention selectively accesses long-range context at arbitrary positions.
@@ 3.1 Model Architecture
As shown in Figure 2(a), RAMba contains Mamba layers, equally divided into upper and lower decoders, akin to previous works [53, 27]. Given the input sequence , their corresponding embeddings are fed into the lower decoder and the outputs are divided into chunks according to chunk size . These chunked representations are then passed through a Transformer-based bi-directional encoder independently to form chunked memories, used for chunk selection and sparse attention in subsequent HSA layers. The upper decoder takes the outputs of the lower decoder as inputs, and alternates between one HSA layer and Mamba layers for times. Between the lower and upper decoders, a chunk selection layer selects the most relevant chunks based on the dot-product similarity between token and chunk representations at each decoding step. HSA layers capture long-range dependencies by attending to the selected chunks.
@@
We focus on architectures like GQA [2], where each group of query heads shares a common KV head, and selects chunks independently. The hidden size is divided into groups, satisfying , with . Here, is the dimension of each group, and is the dimension per head. We denote the -th layer output as , with the -step representation , and the encoded token representations of the -th chunk as . To simplify the notation, we use as a generic symbol to represent one of the groups, as each group behaves in the same way. For example, denotes one of the grouped representations.
@@
At each time step , a token uses the representation derived from the lower decoder to select top- past chunks. Each query group independently selects chunks, which are shared across the heads within the group. Formally, we have:
where is RMSNorm [72], conduct linear transformations, denotes the ranking position in descending order, is the mean-pooled representations of . are representations used for chunk selection, with as their grouped representations. For each group, is the relevance score of to the -th chunk, and is the selected chunk indices for .
@@ 3.1.1 Hierarchical Sparse Attention
To avoid the impact of position encoding on length generalization [29], we discard it and instead model the ordinal relationships of distances akin to the stick-breaking attention [55]. In the stick-breaking process, participants sequentially take a portion from a stick, with later participants dividing the remaining portion left by earlier ones, thus introducing a “most recent” bias. Let the total weights 1 serve as the stick and denote the weight assigned to the -th chunk selected by the -th token. For each group, we have:
@@
where apply linear transform to obtain query, key, and value representations, with query head shares the same KV head. are the key and value corresponding to the -th chunk, while are their grouped representations, shared across all HSA layers. is the query representation of at the -th layer.
The key Difference with MoBA/NSA.
@@ 3.1.2 Kernel Design
In HSA, each token corresponds to a distinct set of chunks, which can lead to a substantial memory footprint in a naive implementation. Inspired by NSA, we address this issue by implementing hardware-aligned HSA kernels based on Triton [57]. As illustrated in Figure 2(b), each GPU thread loads the query representations for a single step along with the selected chunks’ KV pairs for attention computation. Algorithm 1 demonstrates a parallel forward pass for one group of query heads, with multiple groups processed in parallel following the same approach. Particularly, we use softmax-off-by-one [34] to allow tokens in the current chunk to ignore any retrieved tokens. During the back-propagation process, we adopt a two-phased backward pass inspired by Dao [15]. The first stage accumulates gradients for and , followed by and in the second stage. The pseudo-code for this process is shown in Algorithm 2 and 3, in which denotes whether the -th token selects the -th chunk, and represents the index of the -th chunk in .
@@ 3.2 Training & Inference
Many efforts have been made to address Mamba’s length generalization issues. A simple yet effective solution is truncated backpropagation through time (BPTT) [13, 66], in which the first state of a sequence is initialized as the final state of its preceding sequence. We follow this approach in training RAMba and other baselines. However, even though RAMba demonstrates certain extrapolation capabilities, we still observe an increase in perplexity on longer contexts. We hypothesize that RNNs’ memory state provides certain shortcuts [22] for long-range attention, degrading performance on contexts largely exceeding the pre-trained length. Thus we attempt to introduce an appropriate forgetting mechanism into the memory state to disrupt the shortcuts. A straightforward way is memory reset, where sequences are divided into equal segments, and the initial state of each segment is reset to zero. To align with BPTT, we set the initial state as the last hidden state of a random segment in the previous step. In other words, for RNNs, the previous segment is randomly replaced, while the attention mechanism retains access to the original context. Our experiments show that this method effectively improves length generalization for RAMba.
@@
A key challenge in introducing attention mechanisms to RNNs is managing the memory footprint, as the KV cache scales linearly with sequence length. Prior works [35, 27] have demonstrated the feasibility of offloading the KV cache to CPU memory and selectively loading chunks during inference. For a prompt of length , all chunk representations are offloaded to the CPU after prefilling, while only is kept in GPU memory for chunk selection. During decoding, RAMba retrieves and loads chunks at each step for each group. Since the KV cache is shared across all HSA layers, the number of parameters exchanged between the CPU and GPU totals . Our efficiency analysis in § 4.4 demonstrates the overhead for memory exchange is fully acceptable. Since the memory footprint of still increases with the sequence length, to theoretically achieve constant memory, a straightforward way is to offload it to a FAISS [19] database. However, in practice, the memory footprint of is very limited, making such offloading unnecessary. Detailed analysis is elaborated in the experiments section.
@@ 4.1 Setups
To ensure a fair comparison, we pre-train all 370M models from scratch with 4K context length to observe their performance and extrapolation capabilities across various tasks. For 2.7B models, the training details are presented in Appendix E.
@@
We adopt the Mamba-2 architecture as the backbone of the RNN model and YaRN [44] as the Transformer baseline. The parameter size of all models trained from scratch is 370M, with detailed parameters provided in Appendix B. We experiment with Mamba variants with different attention mechanisms, including sliding window attention, native sparse attention (NSA), and HSA. For sliding window attention, the window size is set to 512, incorporating two position encoding schemes: ALiBi [45] and RoPE, the latter following the settings in Samba [51]. We set the chunk size of HSA to 64 following NSA. To ensure that the field of view for sparse attention matches the sliding window size (64 * 8 = 512), we set the number of selected chunks to 8. For NSA, we use its efficient open-source implementation ***https://github.com/fla-org/native-sparse-attention. To isolate the effects of the sparse attention components, we disable the sliding window attention in NSA. The HSA incorporates a single-layered Transformer-based bi-directional encoder for chunk memory encoding, accounting for 5.4% of the total parameters, whose impact on fairness is minimal. HSA layers are inserted into the upper decoder every Mamba layers, with other attention mechanisms like SWA and NSA following the same pattern. These settings remain consistent across all subsequent 370M models. Since the compressed attention in NSA functions similarly to Combiner [50], we do not conduct a separate comparison against Combiner. Some other related works [36] are not included in the experiments due to the lack of open-source implementations.
@@
Ablation Studies.
We denote memory reset usage as w/ m.r. and its absence as w/o m.r.. Additionally, w/o s.b. refers to using softmax without positional encoding instead of stick-breaking weights. For an 8K token context, the memory is reset every 4K tokens, which aligns with the context length of other baselines. However, the chunk selection scope for sparse attention spans 8K tokens, which might be unfair to other baselines. To ensure fair comparisons, we apply the same settings to both HSA and NSA.